OpenStudy (anonymous):

A) the average height of sunflowers in a field is 64 inches with a standard deviation of 3.5 inches. Describe a normal curve for the distribution, including the values on the horizontal axis at one, two, and three standard deviations from the mean B) if there are 3,000 plants in the field, approximately how many will be taller than 71 inches Need Help, last question

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Do you remember the formula for normal curve?

OpenStudy (anonymous):

\[ \Pr(X=x) = \frac{1}{\sigma \sqrt{2\pi}}\exp\left(-\frac{(x-\mu)^2}{2\sigma^2}\right) \]

OpenStudy (anonymous):

in this case \(\exp(x) =e^{x}\), it's notation.

OpenStudy (anonymous):

So \(\mu=64\), the average, and \(\sigma = 3.5\) the std. dev.

OpenStudy (anonymous):

So this means: \[ \Pr(X=x) = \frac{1}{3.5\sqrt{2\pi}}\exp\left(-\frac{(x-64)^2}{2(3.5)^2}\right) \]

OpenStudy (anonymous):

okay I'm still kinda confused how would i solve it from here? @wio

OpenStudy (anonymous):

is x 71?

OpenStudy (anonymous):

@campbell_st could you help me with this?

OpenStudy (campbell_st):

well it looks like  also have a look at this link, that explains it in a simple way http://www.mathsisfun.com/data/standard-normal-distribution.html